OpenStudy (anonymous):

Traveling East How much time is needed for an object traveling with a velocity of 22 m/s eastward to reach a point 242 m to the east?

OpenStudy (austinl):

\(\displaystyle d=rt\) Where, \(d=\text{distance}\\ r=\text{rate}\\ t=\text{time}\) We know two, \(d=242\text{m}\\ r=22\text{m/s}\) So we now have, \(242=22\times t\) How would you solve for t? :)

OpenStudy (austinl):

Do you understand? I have seen you start to respond several times.

OpenStudy (anonymous):

Not quite :/. I'm at the end. 242= 22x t

OpenStudy (austinl):

How would you get t by itself?

OpenStudy (austinl):

Perhaps think of it like this. \(d=rt\) How would you solve for t here?

OpenStudy (anonymous):

Multiply?

OpenStudy (austinl):

It is already a multiplication of r and t. So what do we do to get rid of a multiplication?

OpenStudy (anonymous):

Divide?

OpenStudy (austinl):

Exactly, you would divide each side by r to get t by itself so now we have, \(\displaystyle t=\frac{d}{r}\) Input the values from above, what do you get?

OpenStudy (austinl):

Remember, \(d=242\\ r=22\)

OpenStudy (anonymous):

11 ?

OpenStudy (austinl):

Correct! Good job!

OpenStudy (anonymous):

Thanks!! :D

OpenStudy (austinl):

No problem!